2.5mL of2/5M weak monoacidic base (Kb)=(1)(10)^-12 at 25*C.The concentration of H^+ at equivale points is(Kw=1(10^-14)at25*C
(1)(3.7)(10^-13)M
(2)(3.2)(10^-7)M
(3)(3.2)(10^-2)M
(4)(3.2)(10^-2)M

Grade:

2 Answers

vikas askiitian expert
13 years ago

BOH + H2O  <====>  B+ + OH-

 

1000 ml solution have molarity = 2/5

1ml = 2/5 *10-3

2.5ml = (2/5)*(2.5)*10-3 = 10-3M

now , concentration (OH-) = (kC)1/2

k = 10-12 , C = 10-3 so

[OH-] = [3.33*10-8]

[H+] = 10-14/[OH-] = 3.2*10-7

 

approve if u like my ans

Kushagra Madhukar
3 years ago
Dear student,

The reaction can be written as
BOH + H2O → B+ + OH-
1000 ml solution have molarity = 2/5
1ml = 2/5 *10-3
2.5ml = (2/5)*(2.5)*10-3 = 10-3M
Now , concentration of (OH-) = (kC)1/2
k = 10-12 , C = 10-3
So,
[OH-] = [3.33*10-8]

[H+] = 10-14/[OH-] = 3.2*10-7

Hope it helps.
Thanks and regards,
Kushagra